# Detecting eclipsing double white dwarfs with electromagnetic and gravitational waves

### v1

@@ 3.3 Results
In Table 4, we show the number of eclipsing DWDs detectable by Gaia and LSST in both Model 1 and Model 2. Our results show that 67 and 273 DWDs produced in Model 1 can be detected by Gaia and LSST, respectively. In Model 2, these numbers increase to 186 and 554, respectively. The primary factor contributing to the difference between Model 1 and Model 2 is the choice of the star formation history as described in Section 2, which results in different temperature distributions for the simulated populations of DWDs as shown in Figure 4. Due to the longer period of cooling timescale, the DWDs from Model 1 have overall lower temperature distributions for the primary and companion stars and are therefore less likely to be observed compared to those from Model 2. The distributions of apparent magnitudes for the corresponding DWDs are shown in Figure 5. As we can see, only a small fraction of the DWDs fall within the limiting magnitude, which is consistent with the results reported by Nelemans (2009).
@@
Figure 6 shows the distributions of the main parameters of the DWDs detectable by the Gaia and the LSST-$r$ band for Model 1. According to the distribution of Galactic latitude, about $70\%$ of the DWDs are located within the ±$\pm 20$ degree range of the Galactic plane. In the equatorial coordinate system, compared to the Gaia’s all-sky observations, the DWDs from the LSST observations are mainly located in the southern sky region. From the magnitude distribution, we can see that the magnitude increases with increasing distance. The peak of the magnitude distribution of the DWDs observed by LSST is around 22, which is slightly higher than that of Gaia. The distribution of distances shows that the DWDs detectable by LSST are mainly within 4 kpc and those detectable by Gaia are mainly within 2 kpc. As we can see from Figure 1 and Section 4, the space-borne GW detectors will have a greater distance reach than these optical telescopes, and therefore can detect more DWDs in our Galaxy. For the EM observations, extinction is a critical factor in the calculation of the apparent magnitude. According to Eqs. 1011, extinction is mainly determined by the Galactic latitude, distance, and structure of the Galaxy. The closer the DWDs are to the Galactic plane, the stronger the extinctions. The distance of the DWDs from the observer and the extinction show a proportional linear relationship; the greater the distance from the DWDs, the stronger the extinction.
@@
Figure 7 illustrates the sky distributions of 67 (273) eclipsing DWDs in Model 1 that are detectable by Gaia and LSST. Our results show that LSST can detect 273 eclipsing DWDs in Model 1 and 554 in Model 2. These numbers are lower than those reported by Korol et al. (2019) (hereafter referred to as K19), which are 1100 and 1475 for two different models. Several factors account for this discrepancy. The major factor is the calculation of errors of apparent magnitude of the WDs. K19 effectively uses $\sigma_{\rm{rand}}^{4}$ instead of $\sigma_{\rm{rand}}^{2}$ in Eq. 7, which leads to an underestimation of the photometric errors, especially for fainter WDs. In our validation tests, using $\sigma_{\rm{rand}}^{4}$ for the random photometric errors allows LSST to detect 998 DWDs in Model 1 and 1815 in Model 2, which are more consistent with the numbers reported in K19. In addition, our simulation of the light curves for each LSST bandpass $(ugrizy)$ is based on the number of sampling points (visits) listed in Table 1. In contrast, K19 considers only the LSST-$r$ band and assigns 1000 sampling points to it. The other minor factors include different star formation histories, initial mass functions, and the masses and mass density distributions of the Galactic disk and bulge adopted in our study.
@@ 4.1 GW signals
The GWs emitted by a DWD can be calculated using the quadrupole-moment formula (Peters & Mathews, 1963), which are characterized by two polarizations:
$h_{+}(t)$ $=\mathcal{A}\left(1+\cos^{2}\iota\right)\cos\left(2\pi ft+\pi\dot{f}t^{2}+\phi_{0}+\Phi_{D}(t)\right)\,,$ (13a)
$h_{\times}(t)$ $=2\mathcal{A}\cos\iota\sin\left(2\pi ft+\pi\dot{f}t^{2}+\phi_{0}+\Phi_{D}(t)\right)\,.$ (13b)
Here, the overall amplitude can be written as
$\mathcal{A}=\frac{2(\mathrm{G}\mathcal{M})^{5/3}}{c^{4}d}(\pi f)^{2/3}\,,$ (14)
and the time derivative of GW frequency $f$ is
$\dot{f}=\frac{96}{5}\frac{f}{\mathcal{M}}(\pi f\mathcal{M})^{8/3}\,.$ (15)
The chirp mass $\mathcal{M}\equiv\left(m_{1}m_{2}\right)^{3/5}/\left(m_{1}+m_{2}\right)^{1/5}$, $m_{1}$ and $m_{2}$ are the component masses. The GW phase evolution in Eq. 13 includes initial phase $\phi_{0}$ and Doppler modulation $\Phi_{D}(t)$. The latter can be expressed as follows:
$\Phi_{D}(t)=2\pi f\frac{R_{0}}{c}\sin\theta\cos\left(2\pi f_{m}t-\phi\right)\,,$ (16)
where $R_{0}\approx 1.5\times 10^{8}$ km is the distance between the Earth and the Sun, $f_{m}$= 1/year is the modulation frequency, $(\theta,\phi)$ are the ecliptic coordinates of the source. The GW strain signal measured by a detector can be expressed as:
@@
Here, $F^{+}(t)$ and $F^{\times}(t)$ are the detector’s antenna pattern functions which depend on the source’s location and its orientation relative to the detector. We adopt the expressions of the detector’s antenna pattern functions for TianQin from Hu et al. (2018); Huang et al. (2020), and for LISA from Cornish & Larson (2003). The squared SNR of the signal can be calculated as follows (Finn, 1992)
@@
where $\tilde{S}_{n}\left(f_{0}\right)$ is the single-sided power spectral density (PSD) of detector at GW frequency $f_{0}$, and $T$ is the effective observation time. The PSDs for TianQin and LISA can be found in Hu et al. (2018); Huang et al. (2020) and Robson et al. (2019), respectively.
@@ 4.2 Fisher information matrix
Note that DWDs are typically quasi-monochromatic sources in the mHz band, so not all DWDs in our simulation exhibit a measurable $\dot{f}$ (or $\Sigma_{\dot{f}\dot{f}}$). Here we adopt the commonly used criterion $\dot{f}\,T^{2}\rho>1$ (Adams et al., 2012) to select sources for which we can obtain reliable results from Eqs. 21 and 22. If this criterion is not met, $\dot{f}$ will be excluded from in the parameters to be estimated in our analysis.
@@ 4.3 Results
Table 6 lists the expected numbers of DWDs in Model 1 and Model 2 with $\rho\geq 7(5)$ for TianQin, LISA, and their joint observation. For TianQin, we consider nominal mission duration of 1, 2, and 5 years, which correspond to effective observation times of 0.5, 1, and 2.5 years, respectively. For LISA, we consider nominal mission duration of 1, 2, and 4 years, which correspond to effective observation times of 0.75, 1.5, and 3 years, respectively. For Model 1, TianQin can detect approximately $5\times 10^{3}$ DWDs with SNRs greater than 7 in 2.5 years of effective observation time and LISA can detect approximately $1.7\times 10^{4}$ DWDs with SNRs greater than 7 in 3 years of effective observation time. For Model 2, TianQin can detect approximately $1\times 10^{4}$ DWDs with SNRs greater than 7 in 2.5 years of effective observation time, which is comparable to the results in Huang et al. (2020) based on a DWD population generated from SeBa (Toonen et al., 2017); LISA can detect approximately $3\times 10^{4}$ DWDs with SNRs greater than 7 in 3 years of effective observation time, which is comparable to the results in Korol et al. (2019). The numbers for Model 2 are about twice as many DWDs as the ones for Model 1. This is mainly due to Model 2 generating more DWDs with frequencies in the millihertz. The joint observation of TianQin and LISA can detect an additional 2,000 DWDs compared to LISA alone.
@@
From the top panel of Figure 8, we can see that the frequencies of the DWDs detectable by TianQin and LISA are predominantly in the range of $10^{-3}-10^{-2}$ Hz. The frequency distribution of DWDs detectable by LISA (centered around $2.5\times 10^{-3}$ Hz) exhibits a slightly lower frequency peak compared to TianQin (centered around $3.5\times 10^{-3}$ Hz). This is primarily attributed to the lower sensitivity curve of LISA compared to TianQin for frequencies below $10^{-2}$ Hz, allowing LISA to detect a greater number of lower frequency DWDs. The bottom panel shows the sensitivity curves ($\sqrt{f{\tilde{S}_{n}\left(f\right)}}$) for TianQin and LISA. Among the 16 optical candidates of eclipsing DWDs (marked as stars), we found that 5 of them (marked as red stars) accumulate SNRs greater than 7 (as shown in Table 3) over a nominal 4-year mission duration of LISA.
@@
Figure 9 shows the distributions of parameter estimation accuracy obtained through the FIM analysis for Model 2. Our result yields 3241 (32%), 9760 (29%), and 10356 (29%) DWDs with SNRs between 7 and 10 for TianQin, LISA, and their joint observation, respectively. Only 2% of the DWDs exhibit SNRs greater than 100 for these three detector configurations. In Table 5, we count the cumulative numbers and corresponding percentages of DWDs with parameter estimation relative errors less than 0.1, 0.2, and 0.5 for TianQin, LISA, and their joint observation in Model 2. It can be seen that the total number of relative errors less than 0.1 for each parameter ranges from 1769 to 12218, and the percentage ranges from $12\%$ to $58\%$. In particular, the number of sources with relative errors less than 0.5 is twice as large as the number of sources with relative errors less than 0.1. The relative error for distance is larger compared to the other parameters. From Equation 22 and the distributions of $\Delta{\mathcal{A}}/\mathcal{A}$, $\Delta{f}/f$, and $\Delta{\dot{f}/\dot{f}}$ in Figure 9, we can see that the relative error in $\Delta{\mathcal{D}}/\mathcal{D}$ is mainly contributed by $\Delta{\mathcal{A}}/\mathcal{A}$. About $79\%$ and $76\%$ of the DWDs have relative errors in $\Delta{\mathcal{A}}/\mathcal{A}$ and $\Delta{\mathcal{D}}/\mathcal{D}$ greater than $1$ for TianQin in Model 2. For the sky localization, the numbers of the DWDs with position errors $\Delta{\Omega}<1~{}\mathrm{deg}^{2}$ ($10~{}\mathrm{deg}^{2}$) are 1889 (4974), 4010 (13750), and 4722 (14746) for TianQin, LISA, and their combination, respectively. Given the fields of view of Gaia and LSST are 1.26 $\rm{deg}^{2}$ and 9.6 $\rm{deg}^{2}$, respectively, these sources can be easily targeted by the EM follow-up observations.
@@
In Figure 10, we show the cumulative distributions of the relative errors of chirp mass, distance, and sky location in Model 2. TianQin and LISA can individually detect 216 (543) and 283 (727) DWDs, respectively, with relative errors in $\Delta{\mathcal{M}}/\mathcal{M}$ of less than 0.001 in Model 1 (Model 2). However, for their joint observation, the number reaches 391 (907), an increase of $38\%~{}(24\%)$ over LISA alone in Model 1 (Model 2). For $\Delta{\mathcal{D}}/\mathcal{D}$, TianQin and LISA can individually detect 915 (1769) and 1913 (3931) DWDs, respectively, with relative errors of less than 0.1 in Model 1 (Model 2). For their joint observation, the number reaches 2243 (4599), an increase of $17\%~{}(17\%)$ over LISA alone in Model 1 (Model 2). For $\Delta{\Omega}$, TianQin and LISA can individually detect 926 (1889) and 1896 (4010) DWDs, respectively, with errors of less than 1 deg2 in Model 1 (Model 2). For their joint observation, the number reaches 2275 (4722), an increase of $20\%~{}(18\%)$ over LISA alone in Model 1 (Model 2). Overall, the numbers of the DWDs with errors $\Delta{\mathcal{M}}/\mathcal{M}$, $\Delta{\mathcal{D}}/\mathcal{D}$, and $\Delta{\Omega}$ of less than 0.001, 0.1, and 1 deg2, respectively, have approximately $20\%$ increases in joint observation over LISA alone. Furthermore, combining EM observations will allow a more precise determination of the sky positions of the DWDs, which can facilitate the study of the structure of our Galaxy.
@@
Figure 8: Bottom panel: The characteristic strain of the 15 optically observed eclipsing DWDs (indicated by blue and red stars), along with the characteristic strain of the simulated DWDs (cyan and orange dots) with $\rho>7$ for TianQin and LISA as a function of frequency. The red and blue solid lines represent the sensitivity curves of LISA and TianQin respectively. Top panel: Frequency distributions of simulated DWDs with $\rho>7$.
@@
Figure 10: The cumulative distributions of the relative errors of (a) $\Delta{\mathcal{M}/\mathcal{M}}$, (b) $\Delta{\mathcal{D}}/\mathcal{D}$, and (c) $\Delta{\mathrm{\Omega}}$ for the DWDs detectable by TianQin, LISA, and their joint observation in Model 2.

### v2

@@ 3.3 Results
In Table 3, we show the number of eclipsing DWDs detectable by Gaia and VRO in both Model 1 and Model 2. Our results show that 67 and 273 DWDs produced in Model 1 can be detected by Gaia and VRO, respectively. In Model 2, these numbers increase to 186 and 554, respectively. The primary factor contributing to the difference between Model 1 and Model 2 is the choice of the star formation history as described in Section 2, which results in different temperature distributions for the simulated populations of DWDs as shown in Figure 4. Due to the longer period of cooling timescale, the DWDs from Model 1 have overall lower temperature distributions for the primary and companion stars and are therefore less likely to be observed compared to those from Model 2. The distributions of mean apparent magnitudes for the corresponding DWDs are shown in Figure 5. The mean apparent magnitude refers to the sum of apparent magnitudes at each sampling point on the light curve divided by the number of sampling points (visits). The magnitude at each sampling point takes into account photometric error and extinction. Therefore, although many sources have a mean magnitude below the limiting magnitude, some of their sampling points (especially those during eclipse) exceed the limiting magnitude, and are thus excluded from our analysis. Furthermore, the photometric error bar increases with increasing magnitude, so more higher magnitude sources have a greater chance of being excluded in VRO than in Gaia. As we can see, only a small fraction of the DWDs fall within the limiting magnitude, which is consistent with the results reported by Nelemans (2009).
@@
Figure 6 shows the distributions of the main parameters of the DWDs detectable by the Gaia and the VRO-$r$ band for Model 1. According to the distribution of Galactic latitude, about $70\%$ of the DWDs are located within the ±$\pm 20$ degree range of the Galactic plane. In the equatorial coordinate system, compared to the Gaia’s all-sky observations, the DWDs from the VRO observations are mainly located in the southern sky region. From the magnitude distribution, we can see that the magnitude increases with increasing distance. The peak of the magnitude distribution of the DWDs observed by VRO is around 22, which is slightly higher than that of Gaia. Note that from the distribution shown in the top panel of the third column, the VRO extends to fainter DWDs compared to Gaia. This is mainly because the Gaia $G$-band covers a wider wavelength range of $330~{}\rm{nm}-1050~{}\rm{nm}$, while the VRO-$r$ band covers a range of $560~{}\rm{nm}-706~{}\rm{nm}$. As a result, the Gaia $G$-band collects more photons, resulting in a higher integrated flux. In this work, we used PHOEBE to simulate the flux and converted it to magnitudes. As expected, the flux simulated by PHOEBE in the Gaia $G$-band is higher than in the VRO-$r$ band for the same source. This can already be seen in Fig. 3, where the same source can have a clear difference in magnitude between the Gaia $G$-band and VRO-$r$ band. Therefore, the bright sources in Gaia will appear less bright in VRO, resulting in fewer bright DWDs in the VRO-$r$ band. The distribution of distances shows that the DWDs detectable by VRO are mainly within 4 kpc and those detectable by Gaia are mainly within 2 kpc. As we can see from Figure 1 and Section 4, the space-borne GW detectors will have a greater distance reach than these optical telescopes, and therefore can detect more DWDs in our Galaxy. For the EM observations, extinction is a critical factor in the calculation of the apparent magnitude, which is mainly determined by the Galactic latitude, distance, and structure of the Galaxy. The closer the DWDs are to the Galactic plane, the stronger the extinctions. The distance of the DWDs from the observer and the extinction show a proportional linear relationship; the greater the distance from the DWDs, the stronger the extinction.
@@
Figure 7 illustrates the sky distributions of 67 (273) eclipsing DWDs in Model 1 that are detectable by Gaia and VRO. Our results show that VRO can detect 273 eclipsing DWDs in Model 1 and 554 in Model 2. These numbers are lower than those reported by Korol et al. (2017) (hereafter referred to as K17), which are 1100 and 1475 for two different models. Several factors account for this discrepancy. The major factor is the calculation of errors of apparent magnitude of the WDs. K17 effectively uses $\sigma_{\rm{rand}}^{4}$ instead of $\sigma_{\rm{rand}}^{2}$ in Eq. 3, which leads to an underestimation of the photometric errors, especially for fainter WDs. In our validation tests, using $\sigma_{\rm{rand}}^{4}$ for the random photometric errors allows VRO to detect 998 DWDs in Model 1 and 1815 in Model 2, which are more consistent with the numbers reported in K17. In addition, our simulation of the light curves for each VRO bandpass $(ugrizy)$ is based on the number of sampling points (visits) listed in Table 2. In contrast, K17 considers only the VRO-$r$ band and assigns 1000 sampling points to it. The other minor factors include different star formation histories, initial mass functions, and the masses and mass density distributions of the Galactic disk and bulge adopted in our study.
@@ 4.1 GW signals
The GW strain signal measured by a detector can be expressed as:
@@
Here, $h_{+}(t)$ and $h_{\times}(t)$ are the two polarizations of the GWs emitted by a DWD, which can be calculated using the quadrupole-moment formula (Peters & Mathews, 1963). $F^{+}(t)$ and $F^{\times}(t)$ are the detector’s antenna pattern functions which depend on the source’s location and its orientation relative to the detector. We adopt the expressions of the detector’s antenna pattern functions for TianQin from Hu et al. (2018); Huang et al. (2020), and for LISA from Cornish & Larson (2003). The squared SNR of the signal can be calculated as follows (Finn, 1992)
@@
where $\tilde{S}_{n}\left(f_{0}\right)$ is the single-sided power spectral density (PSD) of detector at GW frequency $f_{0}$, and $T$ is the effective observation time. The PSD for TianQin and LISA can be found in Hu et al. (2018); Huang et al. (2020) and Robson et al. (2019), respectively.
@@ 4.2 Fisher information matrix
Note that DWDs are typically quasi-monochromatic sources in the mHz band, so not all DWDs in our simulation exhibit a measurable $\dot{f}$ (or $\Sigma_{\dot{f}\dot{f}}$). Here we adopt the commonly used criterion $\dot{f}\,T^{2}\rho>1$ (Adams et al., 2012) to select sources for which we can obtain reliable results from Eqs. 9 and 10. If this criterion is not met, $\dot{f}$ will be excluded from in the parameters to be estimated in our analysis.
@@ 4.3 Results
Table 5 lists the expected numbers of DWDs in Model 1 and Model 2 with $\rho\geq 7(5)$ for TianQin, LISA, and their joint observation. For TianQin, we consider nominal mission duration of 1, 2, and 5 years, which correspond to effective observation times of 0.5, 1, and 2.5 years, respectively. For LISA, we consider nominal mission duration of 1, 2, and 4 years, which correspond to effective observation times of 0.75, 1.5, and 3 years, respectively. For Model 1, TianQin can detect approximately $5\times 10^{3}$ DWDs with SNRs greater than 7 in 2.5 years of effective observation time and LISA can detect approximately $1.7\times 10^{4}$ DWDs with SNRs greater than 7 in 3 years of effective observation time. For Model 2, TianQin can detect approximately $1\times 10^{4}$ DWDs with SNRs greater than 7 in 2.5 years of effective observation time, which is comparable to the results in Huang et al. (2020) based on a DWD population generated from SeBa (Toonen et al., 2017); LISA can detect approximately $3\times 10^{4}$ DWDs with SNRs greater than 7 in 3 years of effective observation time, which is comparable to the results in Korol et al. (2017). The numbers of DWDs generated from Model 2 are about twice as many as these from Model 1. This is mainly due to Model 2 generating more DWDs with frequencies in the millihertz. In addition, Model 1 can serve as a lower bound for the models based on astrophysical star formation histories, because a fraction of the DWDs from other models would have already evolved in the burst-like star formation history in Model 1. The joint observation of TianQin and LISA can detect an additional 2,000 DWDs compared to LISA alone.
@@
From the top panel of Figure 8, we can see that the frequencies of the DWDs detectable by TianQin and LISA are predominantly in the range of $10^{-3}-10^{-2}$ Hz. The frequency distribution of DWDs detectable by LISA (centered around $2.5\times 10^{-3}$ Hz) exhibits a slightly lower frequency peak compared to TianQin (centered around $3.5\times 10^{-3}$ Hz). This is primarily attributed to the lower sensitivity curve of LISA compared to TianQin for frequencies below $10^{-2}$ Hz, allowing LISA to detect a greater number of lower frequency DWDs. The bottom panel shows the sensitivity curves ($\sqrt{f{\tilde{S}_{n}\left(f\right)}}$) for TianQin and LISA. Among the 16 optical candidates of eclipsing DWDs (marked as stars), we found that 5 of them (marked as red stars) accumulate SNRs greater than 7 over a nominal 4-year mission duration of LISA.
@@
Figure 9 shows the distributions of parameter estimation accuracy obtained through the FIM analysis for Model 2. Our result yields 3241 (32%), 9760 (29%), and 10356 (29%) DWDs with SNRs between 7 and 10 for TianQin, LISA, and their joint observation, respectively. Only 2% of the DWDs exhibit SNRs greater than 100 for these three detector configurations. In Table 4, we count the cumulative numbers and corresponding percentages of DWDs with parameter estimation relative errors less than 0.1, 0.2, and 0.5 for TianQin, LISA, and their joint observation in Model 2.
The relative error for distance $\Delta{\mathcal{D}}/\mathcal{D}$ is greater in comparison to the other parameters. From Equation 10 and the distributions of $\Delta{\mathcal{A}}/\mathcal{A}$, $\Delta{f}/f$, and $\Delta{\dot{f}/\dot{f}}$ in Figure 9, we can see that the contribution to $\Delta{\mathcal{D}}/\mathcal{D}$ is predominantly attributable to $\Delta{\mathcal{A}}/\mathcal{A}$. In Model 2, approximately $79\%$ and $76\%$ of the DWDs exhibit relative errors in $\Delta{\mathcal{A}}/\mathcal{A}$ and $\Delta{\mathcal{D}}/\mathcal{D}$ exceeding $1$ for TianQin, respectively. For the sky localization, the numbers of the DWDs with position errors $\Delta{\Omega}<1~{}\mathrm{deg}^{2}$ ($10~{}\mathrm{deg}^{2}$) are 1889 (4974), 4010 (13750), and 4722 (14746) for TianQin, LISA, and their combination, respectively. Given the fields of view of Gaia and VRO are 1.26 $\rm{deg}^{2}$ and 9.6 $\rm{deg}^{2}$, respectively, these sources can be easily targeted by the EM follow-up observations.
@@
For Model 1 (Model 2), Table 6 shows the number of DWDs with $\Delta{\mathcal{M}/\mathcal{M}}$, $\Delta{\mathcal{D}}/\mathcal{D}$, and $\Delta{\mathrm{\Omega}}$ of less than 0.001, 0.1, and 1 $\rm{deg^{2}}$ for TianQin, LISA, and their joint observation, respectively. Overall, the joint observation can increase the numbers for LISA by approximately 20%. These sources are likely to be measured with the highest precision by the space-based detectors and can facilitate the subsequent study of the initial mass function and the structure of our Galaxy.
@@
Figure 8: Bottom panel: The characteristic strain of the 15 optically observed eclipsing DWDs (indicated by blue and red stars), along with the characteristic strain of the simulated DWDs (cyan and orange dots) with $\rho>7$ for TianQin and LISA as a function of frequency. The red and blue solid lines represent the sensitivity curves of LISA and TianQin respectively. Top panel: Frequency distributions of simulated DWDs with $\rho>7$. Details on these 15 sources can be found in (Burdge et al., 2020b; Kaplan et al., 2014; Parsons et al., 2011; Vennes et al., 2011; Brown et al., 2011; Kilic et al., 2014; Hallakoun et al., 2016; Brown et al., 2017; Burdge et al., 2020a; Munday et al., 2023; Kosakowski et al., 2023).
@@
Table 6: The number of DWDs with $\Delta{\mathcal{M}/\mathcal{M}}$, $\Delta{\mathcal{D}}/\mathcal{D}$, and $\Delta{\mathrm{\Omega}}$ of less than 0.001, 0.1, and 1 $\rm{deg^{2}}$ for TianQin, LISA and their joint observation in Model 1 (Model 2).
TQ LISA TQ+LISA
$\Delta{\mathcal{M}/\mathcal{M}}<0.001$ 216 (543) 283 (727) 391 (907)
$\Delta{\mathcal{D}/\mathcal{D}}<0.1$ 915 (1769) 1913 (3931) 2243 (4599)
$\Delta{\mathrm{\Omega}}<1\rm{deg^{2}}$ 926 (1889) 1896 (4010) 2275 (4722)